Jill0087:

A certain species of fish has been declining in population by 7%. The population at time zero is 500,000. What is the rate of decay?

Gdeinward:

This is correct right? @Vocaloid

Gdeinward:

Where C is initial number, x is percentage of decrease, and y is the time usually represented by year.

Gdeinward:

\[D = C(1-\frac{ x }{ 100 })^{x}\]

Gdeinward:

so we have D = 500,000(1 - 0.07)^2. The 2 represents 2 years, for example purposes. Can you solve the equation?